Sketch a normal curve for each distribution. Label the x-axis at one, two, and three standard deviations from the mean.
1. Mean = 30; Standard deviation = 5
2. Mean = 95; Standard deviation = 12

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3837)


Problem 1 has been answered previously at this link here
https://www.algebra.com/algebra/homework/Probability-and-statistics/Probability-and-statistics.faq.question.1192102.html

Problem 2 will have the same idea except with these numbers from left to right:
59, 71, 83, 95, 107, 119, 131

95 is at the very center since it's the mean
95-1*12 = 83 and 95+1*12 = 107 are one standard deviation below and above the mean
95-2*12 = 71 and 95+2*12 = 119 are two standard deviations below and above the mean
95-3*12 = 59 and 95+3*12 = 131 are three standard deviations below and above the mean

Answer by ikleyn(53853)

.

In the web-page

https://www.quora.com/How-do-you-plot-a-normal-distribution-in-Excel-with-mean-and-standard-deviation

you will find a complete description on how to plot a normal curve in Excel
with the given mean and the given standard deviation.

To plot a normal distribution in Excel with mean and standard deviation, you can follow these steps:

 1.  Enter the mean and standard deviation values in separate cells.

 2.  In another cell, enter the formula "=NORM.DIST(x,mean,sd,0)" where "x" is the value 
     for which you want to calculate the normal distribution probability, "mean" is the mean value 
     you entered in step 1, "sd" is the standard deviation value you entered in step 1, 
     and "0" indicates that you want the cumulative distribution function.

 3.  Copy the formula from step 2 and paste it into the cells below to calculate the normal distribution probabilities 
     for different values of "x".

 4.  Select the range of cells containing the normal distribution probabilities, including the cell 
     with the formula in step 2.

 5.  Click on the "Insert" tab in the ribbon.

 6.  Click on the "Recommended Charts" button.

 7.  Select the "All Charts" tab.

 8.  Click on the "Statistical" category.

 9.  Select the "Normal Distribution" chart type.

10.  Click "OK" to create the chart.